A mixture consisting of only iron(III) bromide (FeBr3) and aluminum bromide (AlBr3) weighs 0.9553 g. When the mixture is dissolved in water and an excess of silver nitrate is added, all the bromide ions associated with the original mixture are precipitated as insoluble silver bromide (AgBr). The mass of the silver bromide is found to be 1.9182 g. Calculate the mass percentages of iron(III) bromide and aluminum bromide in the original mixture.
1
Expert's answer
2020-09-21T06:38:05-0400

In this answer, use RAM (Fe=55.9, Ag=107.9, Al=27.0, N=14, )=16, Br=79.9)


Step I: Balanced equation for the reaction


AlBr3(aq) + FeBr3(aq) + 6AgNO3(aq) "\to" 6AgBr(s) + Al(NO3)3(aq) + Fe(NO3)3(aq)


Step II: Moles of the product and the reactants


RFM of AgBr = 107.9 + 79.9 = 187.8g/mol


That means 1 mole of AgBr = 187.8g

? moles = 1.9182g

"= \dfrac{1molx1.9182g}{187.8g} = 0.0102moles"


Mole ratio of AgBr:AlBr3 = 6:1


Moles of AlBr3 "=" "\dfrac{1}{6}x0.0102 = 0.0017moles"


Likewise, moles of FeBr3"=\dfrac{1}{6}x0.0102=0.0017moles"


Step III: Mass of the reactants

RFM of AlBr3 = 27 + (3x79.9) = 266.7g/mol


That means 1mol of AlB3 =266.7g

0.0017moles = ?

"=\dfrac{0.0017mol x 266.7g}{1mol}= 0.45339g"



% by mass of AlBr3 in the original mixture"= \dfrac{0.4533g x100}{0.9553g} = 47.46%" %


RFM of FeBr3 = 55.9 + (3 x 79.9) = 295.6g/mol


That means 1 mole of FeBr3=295.6g

0.0017moles = ?


"=\dfrac{0.0017molx 295.6g}{1mol}=0.50252g"


% by mass of FeBr3 in the original sample ="\dfrac{0.50252gx100}{0.9553g}=52.6" %